Ali is 14 years old. His mother is three times as old as he is. In x years' time, his mother will be twice as old as Ali. Form an equation in x and hence find x.
Answer by ikleyn(52797)

.
Ali is 14 years old. His mother is three times as old as he is. In x years' time, his mother will be twice as old as Ali.
Form an equation in x and hence find x.
~~~~~~~~~~~~~~~~~~~~~~~~~

Ali's mother is 3*14 = 42 years now.

In x years . . . 

2*(14 + x) = 42 + x.

28 + 2x = 42 + x,

2x - x = 42 - 28,

x = 14.

Answer. In 14 years.

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

It contains a lot of other word problems and many other interesting and useful things.